OpenStudy (anonymous):

find the values of the indicated trigonometric functions. Find cscθ, given cosθ=0.1063

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

do you have answers I can choose from?

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

mmm ok I think I know but I don't wanna point you in the wrong direction if im wrong

OpenStudy (anonymous):

@Grazes

OpenStudy (anonymous):

Do you know your trig identities?

OpenStudy (anonymous):

1/0.1063=9.41

OpenStudy (anonymous):

not to well no

OpenStudy (anonymous):

Well, you know that csc(θ) = 1/sin(θ), right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Have you learned that sin2 θ + cos2 θ = 1?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

sin^2(θ) + cos^2(θ) = 1 is the function of the unit circle. I don't believe there is a simply way to convert cos to csc, but do you know sin^2(θ) = 1 − cos^2(θ) OR cos^2(θ) = 1 − sin^2(θ)

OpenStudy (jdoe0001):

\(\bf csc(\theta) = \cfrac{1}{sin(\theta)}\qquad if \qquad cos(\theta)=sin(90^o-\theta)\qquad thus\\ \quad \\ csc(\theta) = \cfrac{1}{sin(90^o-\theta)}\)

OpenStudy (jdoe0001):

assuming is 0.1063 degrees, as opposed to radians

OpenStudy (jdoe0001):

\(\bf csc(\theta) = \cfrac{1}{sin(\theta)}\qquad if \qquad cos(\theta)=sin\left(\frac{\pi}{2}-\theta\right)\qquad thus\\ \quad \\ csc(\theta) = \cfrac{1}{sin\left(\frac{\pi}{2}-\theta\right)}\) if it's 0.1063 radians

OpenStudy (anonymous):

1^2+0.1063^2 1=+0.1129969=1.01129969=1.0056

OpenStudy (anonymous):

1/sin 1.00560=56.97

OpenStudy (jdoe0001):

hhh... smokes... I misread.. again hehe... ahemm is just a value..... dohh

OpenStudy (jdoe0001):

lemme redo that

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

\(\bf cos(\theta) = 0.1063\qquad \textit{we know that}\quad cos(\theta) = sin\left(\frac{\pi}{2}-\theta\right)\quad thus\\ \quad \\ cos(\theta) = 0.1063\implies sin\left(\frac{\pi}{2}-\theta\right) = 0.1063\\ \quad \\ \quad \\ csc(\theta) = \cfrac{1}{sin(\theta)}\qquad thus \qquad \cfrac{1}{sin\left(\frac{\pi}{2}-\theta\right)} = \cfrac{1}{0.1063}\)

OpenStudy (jdoe0001):

more or less all we did was, grab the equation \(\bf sin\left(\frac{\pi}{2}-\theta\right) = 0.1063\) and raise both sides at -1

OpenStudy (jdoe0001):

which will give the inverse of sine, which is the cosecant, for the same angle

OpenStudy (anonymous):

ok sol e me see what i get real quick

OpenStudy (anonymous):

0.1061

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{0.1063} \quad ? \)

OpenStudy (anonymous):

oops 9.41

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

lol i just s aw that

OpenStudy (anonymous):

so was doingi t correctly in the first p lace

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

ok great i just wa s unsure

OpenStudy (anonymous):

sorry for the lagg in my computer

OpenStudy (jdoe0001):

tis ok